A water strider sits on a the surface of a still lake. A rock thrown into the lake creates a series of sinusoidal ripples that pass through the location where the strider sits, so that it rises and falls. The distance between its highest and lowest locations is 0.008 m. What is the amplitude of the wave, .002m, .004 m, or .008 m?
1
Expert's answer
2011-04-07T06:44:25-0400
The amplitude is D/2 = 0.008 /2 = 0.004 m.